Cassini Significant Events 12/15/05 - 12/20/05
December 23, 2005
(Source: Cassini Project)
Thursday, December 15 (DOY 349):
The S17 background sequence was sent to the spacecraft today. It is confirmed to be registered and active. The sequence will begin execution this Saturday.
This week the entire suite of Magnetospheric and Plasma Science (MAPS) instruments, which includes the Cassini Plasma Spectrometer (CAPS), Cosmic Dust Analyzer (CDA), Ion and Neutral Mass Spectrometer (INMS), Magnetometer Subsystem (MAG), Magnetospheric Imaging Instrument (MIMI) and Radio and Plasma Wave Science (RPWS), will simultaneously perform low-rate outer magnetospheric surveys to observe the variability of magnetospheric boundaries at a variety of radial distances. Several MAPS instruments will also participate in a campaign to study the interactions between icy satellites, rings, and the magnetosphere.
Optical remote sensing activities include Imaging Science Subsystem (ISS) observations of many small moons for orbit determination, a Visual and Infrared Mapping Spectrometer (VIMS) and Composite Infrared Spectrometer (CIRS) integration of the E Ring near 60 degrees phase to build up spectra of its vertical profile, and a joint ISS and Ultraviolet Imaging Spectrograph (UVIS) observation of Enceladus to obtain longitudinal phase coverage and to measure the ultra-violet albedo across longitudinal phase space.
Friday, December 16 (DOY 350):
As part of the final sequence development process, teams have the opportunity to test any new activities in the Integrated Test Laboratory (ITL). Normally this test is scheduled around week five of this process. For S19, Radio Science requested that we perform the S19 ITL test on the first round of products this week, rather than the later products to take advantage of the extra time around the holidays. Since there were no other items to test, the change of schedule was approved and the test was performed this week.
The Titan Orbiter Science Team hosted a Titan 9 and Titan 10 preview meeting today. The unique aspects of these two flybys are the similarity in geometry to the Voyager flyby, and the first solar occultation we'll observe. The theme of the meeting was, as usual, "what have we learned so far and how does the next set of measurements address our overall science objectives".
Saturday, December 17 (DOY 351):
S17 began execution on-board the spacecraft today. The science instruments were initialized and Instrument Expanded Block files successfully loaded from the SSR. S17 will run for 42 days, completing on January 26, 2006. Two targeted Titan flybys, numbers 9 and 10, occur during this sequence along with non-targeted flybys of Enceladus, Pallene, Telesto, Helene, and Rhea. Orbit Trim Maneuvers (OTM) 46 through 50 are scheduled for this timeframe. OTM-48 was deleted a month or so ago, and it will be reported further on in this report that OTM 46 has been cancelled, leaving three remaining OTMs.
Monday, December 19 (DOY 353):
Navigation, the Spacecraft Operations Office, and the Science Teams met today and were able to officially cancel OTM-46. This reduces the operations activities over the holiday weekend to execution of an Icy Satellite live update on December 24.
An amazingly thin view of Saturn's rings against a backdrop of an enormous Saturn is Astronomy Picture of the Day today.
The Cassini Holiday Party was held today. Flight team members provided an amazing array of desserts as they were entertained the JPL Chorus singing traditional holiday tunes, and by the Cassini Virtual Singers singing "The Twelve Revs of Saturn" and other NON traditional songs (:>). A hit of the event was edible Cassini spacecraft made from ice cream cones, cake and candy provided by Outreach. Here is a link for the edible Cassini: http://www.nasa.gov/lb/audience/forkids/activities/A_Edible_Cassini_Spacecraft.html
Tuesday, December 20 (DOY 354):
Science has requested that operations take a look at a contingency proposal for preserving the Titan 9 data downlinked from the spacecraft on Monday, December 26. This plan would only go into effect if there is a data outage over the DSN pass that day. Being sensitive to asking teams to work on the holiday, the following compromise was reached:
- If there is an outage, no commands will be sent during the Monday evening 70M DSS-14 pass.
- The ACE will page the sequence Leads and the CDS folks during the pass if a DSN outage occurs so they are aware that recovery efforts will need to be done on Tuesday.
- On Tuesday, at the Ops Status meeting, the ACE will report on the extent of any DSN loss, and a small group will determine the best plan to recover the data. In particular we would like UVIS to report at the Ops Status Meeting if they have received the entire HDAC observation data set.
The live update to execute on DOY 358/359 was uplinked to the spacecraft today.
A delivery coordination meeting was held today for Cassini Information Management System(CIMS) software version 3.2.3. Various updates were made to the Resource Checker, a CIMS utility used by Science Planning. The software was installed for operations use right after the meeting.
The most recent spacecraft telemetry was acquired on Tuesday, December 20, from the Goldstone tracking complex. The Cassini spacecraft is in an excellent state of health and all subsystems are operating normally. Check out the Cassini web site at http://saturn.jpl.nasa.gov for the latest press releases and images.